OpenStudy (winsteria):

SOMEONE PLEASE HELP ME IF YOU DO I'LL GIVE YOU A MEDAL AND FAN PLEASE HELP Solve for h. P = mgh A. h = P – mg B. h = pg/m C. h = p/mg D. h = P + mg

OpenStudy (anonymous):

\[30=2\times 3\times 5\] \[5=\frac{30}{2\times 3}\]

OpenStudy (winsteria):

Um

OpenStudy (winsteria):

What do I do? @satellite73

OpenStudy (anonymous):

replace 30 by P, 2 by m, 3 by g

OpenStudy (anonymous):

and 5 by h

OpenStudy (winsteria):

;-;

OpenStudy (anonymous):

it is an analogy if that is confusing, then divide both sides by \(mg\) to solve for \(h\)

OpenStudy (winsteria):

okay

OpenStudy (winsteria):

@cody7744

OpenStudy (winsteria):

B? idk @satellite73

OpenStudy (winsteria):

@Nnesha

Nnesha (nnesha):

mg is multiplying by h so what is opposite of multiply ???

OpenStudy (winsteria):

divide

Nnesha (nnesha):

so divide both side by mg

Nnesha (nnesha):

what you get

OpenStudy (winsteria):

idek

Nnesha (nnesha):

ok so you have to divide both side by mg to get h itself ok

Nnesha (nnesha):

 so you have to divide both side by mg to solve for h now what you should do how you can solve this